OpenStudy (anonymous):

Given the parent function f(x)= log10x, what is the equation of the function shown in the graph?

OpenStudy (anonymous):

the graph points are -3, 2 and 6, 3

OpenStudy (anonymous):

there is a dotted line through -4.

OpenStudy (primeralph):

Just draw it.

OpenStudy (anonymous):

a. f(x)= log10(x - 4) -1 b. f(x)= log10(x+4) +2 c. f(x)= log10(x+4) +4 d. f(x)=log10(x-4) +3

OpenStudy (anonymous):